# Adaptive Attention Distillation for Robust Few-Shot Segmentation under Environmental Perturbations

### v1

@@ VI Conclusion
In this paper, we introduced the Environment-Robust Few-Shot Segmentation (ER-FSS) setting to address a critical gap in existing research: the poor performance of few-shot segmentation models in complex, real-world conditions. To facilitate research in this area, we established the ER-FSS benchmark, a comprehensive collection of datasets designed to evaluate model robustness against environmental challenges like motion blur, camouflage, and viewpoint shifts. We then proposed the Adaptive Attention Distillation (AAD) method, an innovative framework that enhances environmental robustness by learning to distill core, class-discriminative semantics. By iteratively contrasting support and query features to generate a set of compact class queries, AAD effectively guides the model to focus on the essential characteristics of the target category, proving resilient to significant intra-class appearance variations. Extensive experiments demonstrate that our AAD method significantly outperforms existing state-of-the-art FSS and pretrain-finetune approaches on the ER-FSS benchmark, establishing a new baseline for developing practical and robust segmentation models.

### v2

@@ VI Conclusion
In this paper, we introduced the Environment-Robust Few-Shot Segmentation (ER-FSS) setting and benchmark to address the fragility of existing models under real-world perturbations. To tackle these challenges, we proposed the Adaptive Attention Distillation (AAD) framework, which distills robust class semantics via iterative support-query interactions. Despite achieving state-of-the-art robustness, we acknowledge certain limitations. The interactive nature of our module introduces specific computational overhead, and in extremely low-data regimes, such as 1-shot scenarios, our performance occasionally trails behind specialized SOTA methods that are less data-hungry. Future work will therefore focus on optimizing the model’s efficiency and low-shot capability while significantly broadening the ER-FSS benchmark. We aim to incorporate a wider array of scenes and environmental perturbations, alongside other dimensions of difficulty—such as extreme intra-class diversity and acquisition variability—to construct a more holistic and challenging evaluation standard for practical segmentation systems.